# Optimal Design of Automated Market Markers on Decentralized Exchanges

### v2

@@ 1 Introduction
Most studies of AMM on DEX in the literature either isolate LPs’ liquidity provision on DEX from other investment opportunities or implicitly assume that LPs’ investment on CEX has to follow the same allocation as in the liquidity pool on DEX. For example, impermanent loss and loss-versus-rebalancing, which capture the LPs’ losses to arbitrageurs in the liquidity pool and are commonly discussed in the literature, are defined to be the difference between liquidity pool value and the value of the static portfolio with the same initial holding of assets as in the liquidity pool or the replicating portfolio with the same holding of assets as in the liquidity pool at all time; see, for instance, Engel and Herlihy (2022), Angeris and Chitra (2020), Milionis, Moallemi, Roughgarden and Zhang (2024), and Cartea, Drissi and Monga (2023c). Thus, the implicit assumption behind impermanent loss and loss-versus-rebalancing is that the LP compares the liquidity pool on DEX with the same portfolio of assets on CEX. In practice, LPs can hold any portfolio of assets on CEX, while the allocation of assets in the liquidity pool is determined by the pricing function and the exchange rate of assets in the pool, the latter closely following the exchange rate on CEX. Thus, the allocation of assets in the liquidity pool can be inefficient because it can be different from the optimal allocation of assets if the LPs invest on CEX only, e.g., the Merton’s allocation if the LPs maximize the expected utility of their returns. This inefficiency leads to an opportunity cost when LPs provide liquidity on DEX. We are the first in the literature to consider this opportunity cost, and we will show that it plays a critical role in the optimal design of the AMM.
@@ 1.1 Related Literature
The literature has observed that the curvature of the pricing function determines the price slippage and that a higher level of price slippage reduces both LPs’ fee revenue from liquidity traders and LPs’ losses to arbitrageurs. See, for instance, Angeris et al. (2020), Capponi and Jia (2021), and Fabi and Prat (2023). Thus, LPs face a trade-off between high price slippage, which implies a smaller amount of loss to arbitrageurs, and lower price slippage, which implies more fee revenues from liquidity traders. There are a few studies of optimally designing the pricing function in the literature. Capponi and Jia (2021) derive the optimal mixture of the constant sum market maker and the CPMM to maximize the aggregate welfare of LPs, arbitrageurs, investors, and blockchain validators. Lipton and Sepp (2021) use simulation to compute the price slippage and the LP’s profit and loss defined as the fee revenue net the impermanent loss on AMM with a weighted average of the CPMM and the constant sum market maker. They find that both the price slippage and the LP’s profit and loss become higher when the pricing function moves closer to the CPMM. Goyal et al. (2023) derive the optimal pricing function that minimizes the expected inefficiency of AMM, defined as the probability of a trade with a fixed size and tolerance of price slippage not being executed on AMM. Capponi and Chen (2024) find the optimal pricing function that maximizes the expected utility of the LP’s payoff under the same single-period setting as in Capponi and Jia (2021). Milionis et al. (2023b) and Milionis, Moallemi and Roughgarden (2024) describe exchange mechanisms, including DEX with AMM, by demand curves submitted by LPs to exchanges. The authors then study the optimal design of demand curves. Adams et al. (2024) propose a new AMM design, the auction-managed AMM, that can benefit liquidity providers by reducing losses to arbitrageurs and increasing revenue from liquidity traders. The authors show that under certain assumptions, this AMM has higher liquidity in equilibrium than the standard AMM. There are also some studies of providing axiomatic foundations for various pricing functions; see, for instance, Schlegel et al. (2023) and Frongillo et al. (2024).
@@
There are some existing studies on liquidity provisions on DEX with AMM. Chen et al. (2023) consider static liquidity provision and show that the LPs provide liquidity if and only if the exchange rate volatility falls into an interval. Fan et al. (2022) and Heimbach et al. (2022) consider static liquidity provision in Uniswap v3 (Adams et al., 2021). By contrast, we consider dynamic liquidity provision and the optimal design of the liquidity pool. Fan et al. (2023), Bar-On and Mansour (2023), and Cartea, Drissi and Monga (2023a) consider dynamic liquidity provision on Uniswap v3, but they do not consider optimal design of the liquidity pool. Bayraktar et al. (2024) study the competition between multiple LPs through mean-field games.
@@
Angeris et al. (2023) study the portfolio value of the liquidity pool on AMM. The authors show that any concave, nonnegative, nondecreasing, and 1-homogeneous portfolio value function can replicated by some pricing function on AMM. Fabi and Prat (2023) use results in the consumer theory to study the properties of CFMMs. In particular, they show that the cost of trading, captured by the slippage and proportional to the curvature of the bonding curve, is inversely proportional to the divergent loss of the LP. Sabate-Vidales and Šiška (2022) study the optimal trading strategy of an investor on DEX using reinforcement learning algorithms. There are also studies of trading in multiple pools and pool competition; see, for instance, Fritsch (2021), Frtisch et al. (2022), and Zhang et al. (2022). Krishnamachari et al. (2021) and Port and Tiruviluamala (2022) discuss dynamic pricing functions. Cartea et al. (2024) study an optimal liquidation problem for an investor who wants to exchange a large position of an asset for another asset on DEX with AMM in a fixed period of time. The authors consider a continuous-time setting and formulate the problem as a stochastic control problem similar to the optimal liquidation problem on a centralized market (Almgren and Chriss, 2001). Cartea, Drissi and Monga (2023b) extend Cartea et al. (2024) to the case of liquidating multiple assets. Cohen et al. (2023) consider the trading fees gained by the LPs on DEX in a continuous-time model and derive an upper bound for the fees. This upper bound resembles the loss-versus-rebalance proposed by Milionis, Moallemi, Roughgarden and Zhang (2024). Angeris et al. (2022) discuss trading and liquidity provision in multiple-asset pools in a single-period setting. Cartea, Drissi, Sánchez-Betancourt, Siska and Szpruch (2023) propose the so-called arithmetic liquidity pool and study impermanent loss and optimal action of the liquidity provider in this pool. On decentralized lending, leasing, and renting protocol Olympus, by controlling the allocation to bonds and the inflation rate, Chitra et al. (2022) design a liquidity provision framework that improves protocol efficiency.
@@ 2.1 Trading on DEX with AMM
$\underbrace{F(y^{A}-D^{A},y^{B}-D^{B})}_{\text{after the trade}}=\underbrace{F(y^{A},y^{B})}_{\text{before the trade}}.$ (1)
@@
Note that the value of the pricing function $F$ will increase after the trade due to the trading fee; that is,
$\underbrace{F(y^{A}-(1+f\mathbf{1}_{D^{A}<0})D^{A},y^{B}-(1+f\mathbf{1}_{D^{B}<0})D^{B})}_{\text{after the trade}}>F(y^{A}-D^{A},y^{B}-D^{B})=\underbrace{F(y^{A},y^{B})}_{\text{before the trade}}.$
It is worth mentioning that we follow the protocol in Uniswap V2 and assume that the trading fees are deposited in the pool. By contrast, in some existing studies of AMM in the literature, e.g., in Lehar and Parlour (2023) and Milionis, Moallemi, Roughgarden and Zhang (2024), consistent with the protocol in Uniswap v3, the trading fees are assumed to be paid to the LPs directly and thus do not change the value of the pricing function.
@@ 2.2 Liquidity Provision on DEX with AMM
Denote the LP withdrawal (deposit if negative) amounts of asset $A$ and $B$ as $D^{A}$ and $D^{B}$, respectively, such that the subsequent amounts of asset deposits in the pool become $y^{A}-D^{A}$ and $y^{B}-D^{B}$. Consequently, the value of pricing function $F$ also changes along with deposit or withdrawal:
$\underbrace{F(y^{A}-D^{A},y^{B}-D^{B})}_{\text{after LP deposit or withdrawal}}\neq\underbrace{F(y^{A},y^{B})}_{\text{before LP deposit or withdrawal}}.$
In practice, the amount $D^{A}$ and $D^{B}$ may be subject to constraints in various exchanges. A common requirement used in the most popular DEX Uniswap is that the deposit or withdrawal has to keep the ratio between the asset amounts in the pool unchanged, i.e.,
$D^{A}/D^{B}=y^{A}/y^{B}.$
@@ 3.1 The Market
The market participants include a representative liquidity provider and two types of investors: liquidity traders and arbitrageurs. The LP allocates her wealth in asset $A$ and asset $B$ on both the CEX and the DEX, thereby providing liquidity to the DEX, and gains intertemporal utility over consumption. Both the liquidity traders and arbitrageurs maximize their profits on the DEX by exchanging one risky asset for the other, subject to the trading fees. The arbitrageurs and the LP trade based on perfect information about the true exchange rate between the risky assets on CEX, while the liquidity traders trade based on heterogeneous beliefs on the exchange rate, which is a noisy version of the true exchange rate. The detailed objectives of the LP and investors will be described subsequently. Note that liquidity traders are commonly assumed in recent studies of AMM; see, for instance, Milionis, Moallemi, Roughgarden and Zhang (2024), Capponi and Jia (2021), Hasbrouck et al. (2023), and Lehar and Parlour (2023).
@@
the random gross returns of assets $A$ and $B$ due to the shock. The fundamental values change triggers another round of potential trades from the arbitrageurs in Step 5, updating the deposit to $y^{i}_{(k,3)}$. The deposit amounts $y^{i}_{(k,3)}$ and the post-shock fundamental values $p^{i}_{k+1}$ are carried forward as $y^{i}_{k+1}$ and $p^{i}_{k+1}$ at the end of the $(k+1)$-th period, which is also the beginning of the $(k+2)$-th period. Steps 3 and 5 in our model are, respectively, the reversion arbitrage and the rebalancing arbitrage in Milionis, Moallemi, Roughgarden and Zhang (2024).
@@
    $F_{x}(x,y)/F_{y}(x,y)=G(x/y)$ for some continuously differentiable univariate function $G$ on $(0,\infty)$ satisfying
    $G^{\prime}(z)<0,\;\forall z>0,\quad\lim_{z\downarrow 0}G(z)=\infty,\quad\lim_{z\uparrow\infty}G(z)=0.$
@@
With Assumption 3.1-(i), for any fixed deposits in the pool $y^{A}>0$ and $y^{B}>0$, the average exchange rate of acquiring asset $A$ for a small amount $D_{A}>0$ is well defined, which is $-D_{B}/D_{A}$ with $D_{A}$ and $D_{B}$ satisfying (1). We define the marginal exchange rate of acquiring asset $A$ to be the average exchange rate when the amount $D_{A}$ of asset $A$ to be acquired is infinitesimally small. Applying the implicit function theorem, we can derive from (1) that the marginal exchange rate of acquiring asset $A$ is
$F_{x}(y^{A},y^{B})/F_{y}(y^{A},y^{B}).$
Similarly, we can define the average exchange rate of acquiring asset $B$ for a small amount and the marginal exchange rate of acquiring asset $B$, and the latter is $F_{y}(y^{A},y^{B})/F_{x}(y^{A},y^{B})$, which is the reciprocal of the marginal exchange rate of acquiring asset $A$.
@@
Note that a monotone transformation of the pricing function does not change the exchange rate on AMM. Indeed, consider any strictly increasing, continuously differentiable function $g$. Define the new pricing function $\tilde{F}(x,y):=g(F(x,y))$. Then, the pricing equation (1) holds if and only if
$\tilde{F}(y^{A}-D^{A},y^{B}-D^{B})=\tilde{F}(y^{A},y^{B}).$
It is straightforward to verify that Assumption 3.1 holds for $F$ if and only if it holds of $\tilde{F}$.
@@ 3.2 Investors’ Problem
where
$\phi^{i}(\alpha,\beta):=1-(1+f\mathbf{1}_{\varphi^{i}(\alpha,\beta)<0})\varphi^{i}(\alpha,\beta),\quad i\in\{A,B\}.$ (13)
@@
The ratio of the marginal exchange rate on AMM and the investor’s belief of the exchange rate after the trade (post-trade ratio) depends on the ratio before the trade (pre-trade ratio) as follows
$\underbrace{G\left(\frac{y^{A}-(1+f\mathbf{1}_{\hat{D}^{A}<0})\hat{D}^{A}}{y^{B}-(1+f\mathbf{1}_{\hat{D}^{B}<0})\hat{D}^{B}}\right)/(a/b)}_{\text{post-trade ratio}}=H\left(\underbrace{G\left(\frac{y^{A}}{y^{B}}\right)/(a/b)}_{\text{pre-trade ratio}},a/b\right),$ (14)
@@
Note that $H(s,z)\in[1/(1+f),1+f]$ for all $s>0$ and $z>0$, i.e., the post-trade ratio lies in $[1/(1+f),1+f]$; otherwise the trading will continue. If the pre-trade ratio is in the region $[1/(1+f),1+f]$, there is no trade, so the post-trade ratio remains the same. For $s<1/(1+f)$, by Proposition 3.1-(i), $\varphi^{B}\left(z,G^{-1}\big{(}sz\big{)}\right)<0$, so $H(s,z)>1/(1+f)$. Thus, when the pre-trade ratio is strictly less than $1/(1+f)$, the investor exchanges for asset $A$ and the trading brings the ratio not onto the boundary of the no-trading region $[1/(1+f),1+f]$ but into the interior of the no-trading region. This is because the trading cost is deposited into the liquidity pool and thus changes the exchange rate on AMM. Moreover, by Proposition 3.1-(i), $\varphi^{B}\left(z,G^{-1}\big{(}sz\big{)}\right)$ is strictly increasing in $s$ for $s<1/(1+f)$ and thus $H(s,z)$ is strictly decreasing in $s$ for $s<1/(1+f)$. Consequently, the post-trade ratio is strictly decreasing in the pre-trade ratio on the left side of the no-trading region. This is because the further the pre-trade ratio is away from the no-trading region from the left side, the more valuable asset $A$ is from the investor’s perspective and thus a larger amount of asset $A$ is exchanged by the investor, which incurs a larger trading cost deposited in the liquidity pool and pushes the post-trade ratio further into the no-trading region. Similarly, when the pre-trade ratio is strictly larger than $1+f$, trading occurs and brings the post-trade ratio into the interior of the no-trading region, and the latter is strictly decreasing in the former. See Figure 2 for an illustration of the post-trade ratio versus the pre-trade ratio.
Figure 2: Post-trade ratio with respect to the pre-trade ratio, i.e., the plot of $H(s,z)$ as in (15) with respect to $s$. We use the CGMMM pricing function (2), in which case $H(s,z)$ is independent of $z$; see Section 3.4.3. In the plot, we set $f=1\%$ and $\eta=0.5$, and set other parameter values as listed in Table 1.

### v3

# Optimal Design of Automated Market Makers on Decentralized Exchanges

@@ Abstract
Keywords: automated market maker, decentralized exchange, dynamic programming
@@ 1 Introduction
Most studies of AMM on DEX in the literature either isolate LPs’ liquidity provision on DEX from other investment opportunities or implicitly assume that LPs’ investment on CEX has to follow the same allocation as in the liquidity pool on DEX. For example, impermanent loss and loss-versus-rebalancing, which capture the LPs’ losses to arbitrageurs in the liquidity pool and are commonly discussed in the literature, are defined to be the difference between liquidity pool value and the value of the static portfolio with the same initial holding of assets as in the liquidity pool or the replicating portfolio with the same holding of assets as in the liquidity pool at all time; see, for instance, Engel and Herlihy (2022), Angeris and Chitra (2020), Milionis et al. (2024b), and Cartea et al. (2023c). Thus, the implicit assumption behind impermanent loss and loss-versus-rebalancing is that the LP compares the liquidity pool on DEX with the same portfolio of assets on CEX. In practice, LPs can hold any portfolio of assets on CEX, while the allocation of assets in the liquidity pool is determined by the pricing function and the exchange rate of assets in the pool, the latter closely following the exchange rate on CEX. Thus, the allocation of assets in the liquidity pool can be inefficient because it can be different from the optimal allocation of assets if the LPs invest on CEX only, e.g., the Merton’s allocation if the LPs maximize the expected utility of their returns. This inefficiency leads to an opportunity cost when LPs provide liquidity on DEX. We are the first in the literature to consider this opportunity cost, and we will show that it plays a critical role in the optimal design of the AMM.
@@ 1.1 Related Literature
The literature has observed that the curvature of the pricing function determines the price slippage and that a higher level of price slippage reduces both LPs’ fee revenue from liquidity traders and LPs’ losses to arbitrageurs. See, for instance, Angeris et al. (2020), Capponi and Jia (2021), and Fabi and Prat (2023). Thus, LPs face a trade-off between high price slippage, which implies a smaller amount of loss to arbitrageurs, and lower price slippage, which implies more fee revenues from liquidity traders. There are a few studies of optimally designing the pricing function in the literature. Capponi and Jia (2021) derive the optimal mixture of the constant sum market maker and the CPMM to maximize the aggregate welfare of LPs, arbitrageurs, investors, and blockchain validators. Lipton and Sepp (2021) use simulation to compute the price slippage and the LP’s profit and loss defined as the fee revenue net the impermanent loss on AMM with a weighted average of the CPMM and the constant sum market maker. They find that both the price slippage and the LP’s profit and loss become higher when the pricing function moves closer to the CPMM. Goyal et al. (2023) derive the optimal pricing function that minimizes the expected inefficiency of AMM, defined as the probability of a trade with a fixed size and tolerance of price slippage not being executed on AMM. Capponi and Chen (2024) find the optimal pricing function that maximizes the expected utility of the LP’s payoff under the same single-period setting as in Capponi and Jia (2021). Milionis et al. (2023b) and Milionis et al. (2024a) describe exchange mechanisms, including DEX with AMM, by demand curves submitted by LPs to exchanges. The authors then study the optimal design of demand curves. Adams et al. (2024) propose a new AMM design, the auction-managed AMM, that can benefit liquidity providers by reducing losses to arbitrageurs and increasing revenue from liquidity traders. The authors show that under certain assumptions, this AMM has higher liquidity in equilibrium than the standard AMM. There are also some studies of providing axiomatic foundations for various pricing functions; see, for instance, Schlegel et al. (2023) and Frongillo et al. (2024).
@@
There are some existing studies on liquidity provisions on DEX with AMM. Chen et al. (2023) consider static liquidity provision and show that the LPs provide liquidity if and only if the exchange rate volatility falls into an interval. Fan et al. (2022) and Heimbach et al. (2022) consider static liquidity provision in Uniswap v3 (Adams et al., 2021). By contrast, we consider dynamic liquidity provision and the optimal design of the liquidity pool. Fan et al. (2023), Bar-On and Mansour (2023), and Cartea et al. (2023a) consider dynamic liquidity provision on Uniswap v3, but they do not consider optimal design of the liquidity pool. Bayraktar et al. (2024) study the competition between multiple LPs through mean-field games.
@@
Angeris et al. (2023) study the portfolio value of the liquidity pool on AMM. The authors show that any concave, nonnegative, nondecreasing, and 1-homogeneous portfolio value function can replicated by some pricing function on AMM. Fabi and Prat (2023) use results in the consumer theory to study the properties of CFMMs. In particular, they show that the cost of trading, captured by the slippage and proportional to the curvature of the bonding curve, is inversely proportional to the divergent loss of the LP. Sabate-Vidales and Šiška (2022) study the optimal trading strategy of an investor on DEX using reinforcement learning algorithms. There are also studies of trading in multiple pools and pool competition; see, for instance, Fritsch (2021), Frtisch et al. (2022), and Zhang et al. (2022). Krishnamachari et al. (2021) and Port and Tiruviluamala (2022) discuss dynamic pricing functions. Cartea et al. (2024) study an optimal liquidation problem for an investor who wants to exchange a large position of an asset for another asset on DEX with AMM in a fixed period of time. The authors consider a continuous-time setting and formulate the problem as a stochastic control problem similar to the optimal liquidation problem on a centralized market (Almgren and Chriss, 2001). Cartea et al. (2023b) extend Cartea et al. (2024) to the case of liquidating multiple assets. Cohen et al. (2023) consider the trading fees gained by the LPs on DEX in a continuous-time model and derive an upper bound for the fees. This upper bound resembles the loss-versus-rebalance proposed by Milionis et al. (2024b). Angeris et al. (2022) discuss trading and liquidity provision in multiple-asset pools in a single-period setting. Cartea et al. (2023d) propose the so-called arithmetic liquidity pool and study impermanent loss and optimal action of the liquidity provider in this pool. On decentralized lending, leasing, and renting protocol Olympus, by controlling the allocation to bonds and the inflation rate, Chitra et al. (2022) design a liquidity provision framework that improves protocol efficiency.
@@ 2.1 Trading on DEX with AMM
$F(y^{A}-D^{A},y^{B}-D^{B})=F(y^{A},y^{B}).$ (1)
@@
Note that the value of the pricing function $F$ will increase after the trade due to the trading fee; that is, $F(y^{A}-(1+f\mathbf{1}_{D^{A}<0})D^{A},y^{B}-(1+f\mathbf{1}_{D^{B}<0})D^{B})>F(y^{A}-D^{A},y^{B}-D^{B})=F(y^{A},y^{B}).$ It is worth mentioning that we follow the protocol in Uniswap V2 and assume that the trading fees are deposited in the pool. By contrast, in some existing studies of AMM in the literature, e.g., in Lehar and Parlour (2023) and Milionis et al. (2024b), consistent with the protocol in Uniswap v3, the trading fees are assumed to be paid to the LPs directly and thus do not change the value of the pricing function.
@@ 2.2 Liquidity Provision on DEX with AMM
Denote the LP withdrawal (deposit if negative) amounts of asset $A$ and $B$ as $D^{A}$ and $D^{B}$, respectively, such that the subsequent amounts of asset deposits in the pool become $y^{A}-D^{A}$ and $y^{B}-D^{B}$. Consequently, the value of pricing function $F$ also changes along with deposit or withdrawal: $F(y^{A}-D^{A},y^{B}-D^{B})\neq F(y^{A},y^{B}).$ In practice, the amount $D^{A}$ and $D^{B}$ may be subject to constraints in various exchanges. A common requirement used in the most popular DEX Uniswap is that the deposit or withdrawal has to keep the ratio between the asset amounts in the pool unchanged, i.e., $D^{A}/D^{B}=y^{A}/y^{B}.$
@@ 3.1 The Market
The market participants include a representative liquidity provider and two types of investors: liquidity traders and arbitrageurs. The LP allocates her wealth in asset $A$ and asset $B$ on both the CEX and the DEX, thereby providing liquidity to the DEX, and gains intertemporal utility over consumption. Both the liquidity traders and arbitrageurs maximize their profits on the DEX by exchanging one risky asset for the other, subject to the trading fees. The arbitrageurs and the LP trade based on perfect information about the true exchange rate between the risky assets on CEX, while the liquidity traders trade based on heterogeneous beliefs on the exchange rate, which is a noisy version of the true exchange rate. The detailed objectives of the LP and investors will be described subsequently. Note that liquidity traders are commonly assumed in recent studies of AMM; see, for instance, Milionis et al. (2024b), Capponi and Jia (2021), Hasbrouck et al. (2023), and Lehar and Parlour (2023).
@@
the random gross returns of assets $A$ and $B$ due to the shock. The fundamental values change triggers another round of potential trades from the arbitrageurs in Step 5, updating the deposit to $y^{i}_{(k,3)}$. The deposit amounts $y^{i}_{(k,3)}$ and the post-shock fundamental values $p^{i}_{k+1}$ are carried forward as $y^{i}_{k+1}$ and $p^{i}_{k+1}$ at the end of the $(k+1)$-th period, which is also the beginning of the $(k+2)$-th period. Steps 3 and 5 in our model are, respectively, the reversion arbitrage and the rebalancing arbitrage in Milionis et al. (2024b).
@@
    $F_{x}(x,y)/F_{y}(x,y)=G(x/y)$ for some continuously differentiable univariate function $G$ on $(0,\infty)$ satisfying $G^{\prime}(z)<0,\;\forall z>0,~{}~{}\lim_{z\downarrow 0}G(z)=\infty,~{}~{}\lim_{z\uparrow\infty}G(z)=0.$
@@
With Assumption 3.1-(i), for any fixed deposits in the pool $y^{A}>0$ and $y^{B}>0$, the average exchange rate of acquiring asset $A$ for a small amount $D_{A}>0$ is well defined, which is $-D_{B}/D_{A}$ with $D_{A}$ and $D_{B}$ satisfying (1). We define the marginal exchange rate of acquiring asset $A$ to be the average exchange rate when the amount $D_{A}$ of asset $A$ to be acquired is infinitesimally small. Applying the implicit function theorem, we can derive from (1) that the marginal exchange rate of acquiring asset $A$ is $F_{x}(y^{A},y^{B})/F_{y}(y^{A},y^{B}).$ Similarly, we can define the average exchange rate of acquiring asset $B$ for a small amount and the marginal exchange rate of acquiring asset $B$, and the latter is $F_{y}(y^{A},y^{B})/F_{x}(y^{A},y^{B})$, which is the reciprocal of the marginal exchange rate of acquiring asset $A$.
@@
Note that a monotone transformation of the pricing function does not change the exchange rate on AMM. Indeed, consider any strictly increasing, continuously differentiable function $g$. Define the new pricing function $\tilde{F}(x,y):=g(F(x,y))$. Then, the pricing equation (1) holds if and only if $\tilde{F}(y^{A}-D^{A},y^{B}-D^{B})=\tilde{F}(y^{A},y^{B}).$ It is straightforward to verify that Assumption 3.1 holds for $F$ if and only if it holds of $\tilde{F}$.
@@ 3.2 Investors’ Problem
where $\text{where }\phi^{i}(\alpha,\beta)$ $:=1-(1+f\mathbf{1}_{\varphi^{i}(\alpha,\beta)<0})\varphi^{i}(\alpha,\beta).$ (13)
@@
The ratio of the marginal exchange rate on AMM and the investor’s belief of the exchange rate after the trade (post-trade ratio, left hand side of (14)) depends on the ratio before the trade (pre-trade ratio, first argument of $H$ on right hand side of (14)) as follows
$G\left(\frac{y^{A}-(1+f\mathbf{1}_{\hat{D}^{A}<0})\hat{D}^{A}}{y^{B}-(1+f\mathbf{1}_{\hat{D}^{B}<0})\hat{D}^{B}}\right)/(a/b)=H\left(G\left(\frac{y^{A}}{y^{B}}\right)/(a/b),a/b\right),$ (14)
@@
Note that $H(s,z)\in[1/(1+f),1+f]$ for all $s>0$ and $z>0$, i.e., the post-trade ratio lies in $[1/(1+f),1+f]$; otherwise the trading will continue. If the pre-trade ratio is in the region $[1/(1+f),1+f]$, there is no trade, so the post-trade ratio remains the same. For $s<1/(1+f)$, by Proposition 3.1-(i), $\varphi^{B}\left(z,G^{-1}\big{(}sz\big{)}\right)<0$, so $H(s,z)>1/(1+f)$. Thus, when the pre-trade ratio is strictly less than $1/(1+f)$, the investor exchanges for asset $A$ and the trading brings the ratio not onto the boundary of the no-trading region $[1/(1+f),1+f]$ but into the interior of the no-trading region. This is because the trading cost is deposited into the liquidity pool and thus changes the exchange rate on AMM. Moreover, by Proposition 3.1-(i), $\varphi^{B}\left(z,G^{-1}\big{(}sz\big{)}\right)$ is strictly increasing in $s$ for $s<1/(1+f)$ and thus $H(s,z)$ is strictly decreasing in $s$ for $s<1/(1+f)$. Consequently, the post-trade ratio is strictly decreasing in the pre-trade ratio on the left side of the no-trading region. This is because the further the pre-trade ratio is away from the no-trading region from the left side, the more valuable asset $A$ is from the investor’s perspective and thus a larger amount of asset $A$ is exchanged by the investor, which incurs a larger trading cost deposited in the liquidity pool and pushes the post-trade ratio further into the no-trading region. Similarly, when the pre-trade ratio is strictly larger than $1+f$, trading occurs and brings the post-trade ratio into the interior of the no-trading region, and the latter is strictly decreasing in the former.